# CoCoEvo: Co-Evolution of Programs and Test Cases to Enhance Code Generation

### v1

@@ Abstract
Large Language Models (LLMs) have shown remarkable performance in automated code generation. However, existing approaches often rely heavily on pre-defined test cases, which become impractical in scenarios where such cases are unavailable. While prior works explore filtering techniques between programs and test cases, they overlook the refinement of test cases. To address this limitation, we introduce CoCoEvo, a novel LLM-based co-evolution framework that simultaneously evolves programs and test cases. CoCoEvo eliminates the dependency on pre-defined test cases by generating both programs and test cases directly from natural language problem descriptions and function headers. The framework employs specialized evolutionary operators, including LLM-based crossover and mutation operators for program evolution, along with a test case generation operator for test case evolution. Additionally, we propose optimization strategies such as a crossover rate scheduler to balance exploration and convergence, and a multi-objective optimization method for test case selection. Experimental results on multiple state-of-the-art LLMs demonstrate that CoCoEvo surpasses existing methods, achieving state-of-the-art performance in automated code generation and testing. These results underscore the potential of co-evolutionary techniques in advancing the field of automated programming.
@@
Despite their impressive performance, LLMs do not always generate accurate or correct programs. To enhance the reliability of program generation, test cases are often utilized to evaluate code correctness and provide feedback. Existing approaches frequently rely on pre-defined test cases for evaluation. For instance, Sampling+Filtering [5] first generates a batch of candidate codes, evaluates them using pre-defined test cases, and selects the code with the highest score as the final solution. Other methods, such as Reflexion [6], Self-Repair [7], INTERVENOR [8], LDB [9], and PairCoder [10], utilize execution feedback on pre-defined test cases to refine or repair generated code. These methods generally assume that pre-defined test cases are accurate and can be fully trusted.
However, pre-defined test cases are not always available. In test-driven development (TDD) [11], test cases are defined before program implementation. In real-world scenarios, software development frequently relies on natural language requirements, with no pre-existing test cases or example programs. Furthermore, creating pre-defined test cases can be labor-intensive, requiring significant manual effort. In situations where LLMs must generate test cases independently, the effectiveness of existing methods remains uncertain.
Some approaches do not rely on pre-defined test cases. For instance, methods like CodeT [12] and MBR-Exec [13] generate both code and test cases using LLMs and filter the code based on the execution results of these test cases. Similarly, CodeCoT [14], which utilizes chain-of-thought prompting [15], enables LLMs to generate code and test cases simultaneously and evaluate and refine the code based on these cases. Additionally, AgentCoder [16] introduces a multi-agent system with a dedicated Tester agent to generate test cases for evaluating and repairing code. However, studies such as [12, 17] highlight that LLM-generated test cases can contain significant errors, potentially leading to unreliable evaluations and incorrect feedback. Moreover, many of these methods neglect to verify the quality of the generated test cases. While approaches like CodeT and MBR-Exec implement basic filtering mechanisms for test cases, other methods often use these generated test cases directly for code evaluation without ensuring their correctness.
@@
To address these limitations, we propose CoCoEvo, an LLM-based co-evolution framework that enables the simultaneous evolution of programs and test cases to produce more accurate outcomes (illustrated in Fig. 1). In CoCoEvo, both test cases and programs are generated by LLMs, requiring only natural language problem descriptions and function headers. No pre-defined programs, test cases, or templates are needed. The CoCoEvo framework consists of two alternating steps: program evolution and test case evolution. In program evolution, we introduce LLM-based crossover and mutation operators to generate program offspring. The current population of test cases is used to evaluate program fitness. In test case evolution, we design an additional test case generation operator to produce test case offspring. Conversely, the current population of programs is used to evaluate test case fitness. For programs, we implement a fitness calculation method inspired by CodeT [12], which assesses agreement between programs and test cases. For test cases, we adopt a multi-objective optimization approach to select those with high accuracy and strong discriminatory power, using the Pareto method [22]. Additionally, we introduce a dynamic crossover rate scheduling mechanism, enabling broader exploration during early evolution stages and faster convergence later.
To evaluate the effectiveness of CoCoEvo, we collected a LeetCode-Contest dataset comprising relatively new programming problems that are unlikely to have been included in the training data of the evaluated LLMs. We conducted experiments on four state-of-the-art models: GPT-4o-mini [1], Qwen2.5-Coder-32B [3], Llama-3.1-70B [2], and DeepSeek-V3 [4]. We compared with both basic code generation methods such as Sampling [5] and Sampling+Filtering [5], repair-based methods, and agreement-based methods. The results demonstrate that CoCoEvo outperforms all existing methods. Additionally, the experiments revealed that methods relying on pre-defined test cases for feedback experience significant performance degradation when adapted to self-generated test cases. This highlights the limitations of such methods in scenarios lacking pre-defined test cases. Furthermore, we conducted in-depth comparative experiments and discussions, which validated the effectiveness of the optimization strategies employed in CoCoEvo.
@@
Optimized Evolution Techniques. For program evolution, we design a dynamic crossover rate scheduler that enhances exploration early on and accelerates convergence later. For test cases, we develop a multi-objective optimization method for fitness evaluation and selection.
@@
Experimental Validation. We introduce the LeetCode-Contest dataset and conduct experiments using four state-of-the-art LLMs: GPT-4o-mini, Qwen2.5-Coder-32b, Llama-3.1-70B, and DeepSeek-V3. The experimental results demonstrate that CoCoEvo achieves state-of-the-art performance, surpassing existing methods. Additionally, we conduct ablation studies to evaluate the utility of each module within the framework, further highlighting the effectiveness of our approach.
The remainder of this article is organized as follows. Section II discusses the related work. In Section III, the design details and implementation methods of the CoCoEvo are described. Sections IV and V present the experimental settings and discussions, respectively. Finally, Section VI concludes this article and outlines future work.
@@ II-A Enhancing Code Generation Accuracy with Test Case
The integration of test cases has emerged as a pivotal strategy for improving the quality of code generated by LLMs. Numerous studies leverage test cases to filter, evaluate, and refine generated programs, leading to higher-performing and more reliable code. Below, we provide an overview of notable methodologies in this domain.
Li et al. [5] introduce a Sampling+Filtering approach, where a large set of candidate programs is sampled and scored based on performance against a provided test set. Chen et al. [12] propose a framework in which both code and test cases are generated by LLMs simultaneously, using a dual agreement metric to identify and select better programs. Shi et al. [13] assess semantic similarity by executing generated programs on test cases and leveraging the execution results to guide program selection.
Beyond evaluation, test cases play a critical role in iterative repair and refinement of generated code. Execution outcomes from test cases can be used to repair faulty programs. For instance, Shinn et al. [6] introduce the Reflexion framework, where failure messages from test executions are converted into verbal feedback for program repair. Similarly, Olausson et al. [7] evaluate LLMs’ capabilities to repair programs using failed test feedback. Zhang et al. [23] allow LLMs to analyze test execution results to revise erroneous code. Zhong et al. [9] extend this idea by enabling step-by-step debugging. Huang et al. [14] adopt a chain-of-thought prompting [15] approach that generates code and tests sequentially, iteratively improving the code based on test feedback. Zhang et al. [10] introduce a pair programming framework where test feedback informs repair strategies and planning.
The generation of test cases differs across these methodologies. For example, CodeT [12] and CodeCoT [14] rely entirely on LLMs to generate test cases, whereas Reflexion employs LLMs primarily for simple programming tasks. Most other methods, however, assume pre-defined test cases as part of their evaluation frameworks, leveraging them for filtering, refinement, and program selection.
@@ II-B Automated Test Case Generation
Early works, such as [24], leveraged search-based methods to generate test cases. More recently, there has been a growing trend toward utilizing LLMs for test case generation. Studies such as [25], [26], and [27] have explored the effectiveness of LLMs in automating program testing. Similarly, [28] investigates methods for enhancing existing test cases using LLMs. Other approaches, such as TOGA [29], TOGLL [30], and the method described in [31], focus on generating test oracles based on predefined templates. Meanwhile, [32] introduces a model trained with deep reinforcement learning for text-to-test case generation. The application of LLMs in software development extends to generating test cases for various scenarios. For instance, [33] employs GPT-4 to generate security tests for identifying vulnerabilities, while [27] uses LLMs to assist in generating unit tests for automatic software testing. Similarly, [34] demonstrates the use of LLMs to create test cases aimed at reproducing general software bugs. LLMs are also integrated into code generation studies to improve the accuracy of generated code through test case generation. CodeT [12], for example, utilizes LLMs with zero-shot prompts to directly generate test cases. Reflexion [6] introduces a Test Agent powered by LLMs for test case generation.
@@ II-C Evolution with LLMs for Code Generation
The integration of LLMs with evolutionary algorithms has been extensively explored to enhance code generation. For example, Bradley et al. [35] introduced OpenELM, an open-source Python library that designs specialized evolutionary operators for code generation. Chen et al. [36] utilized LLMs as adaptive mutation and crossover operators in evolutionary neural architecture search (NAS), optimizing neural network design by merging LLM capabilities with evolutionary processes. Liu et al. [37] proposed an automated approach for evolving optimization algorithms using LLMs, which eliminates the need for manual design by applying standard evolutionary operations such as initialization, selection, and mutation. In a related study, Liu et al. [38] developed a framework for evolving heuristics by integrating LLMs with evolutionary computation to co-evolve both natural language descriptions and executable code for heuristic design, further extending previous methodologies.
Ma et al. [39] introduced an algorithm combining LLMs with evolutionary search to generate and optimize reward functions for reinforcement learning tasks, incorporating a reward reflection mechanism for refinement. Meyerson et al. [40] demonstrated the use of few-shot prompting with LLMs to generate offspring in evolutionary algorithms, enabling a general crossover operator by concatenating parent solutions into a single prompt, which can adapt to various domains. Similarly, Romera et al. [41] integrated LLMs with evolutionary algorithms to create context-aware program mutations, eliminating the need for manually defined mutation operators. Ye et al. [42] proposed ReEvo, which combines evolutionary search with LLM-driven reflections to generate optimization heuristics, efficiently exploring heuristic spaces while offering verbal gradients for optimization. Lastly, AutoTest [43] combines automated test case generation with code execution in an evolutionary framework, utilizing LLMs to co-generate code solutions and their corresponding test cases for enhanced optimization.
Additionally, various LLM-based co-evolution methods have been proposed. For instance, Chi et al. [19] identifies outdated test cases by co-evolving production and test code. Ruan et al. [20] constructs test oracles from bug reports using predefined templates. Chi et al. [19] leveraged LLMs with dynamic validation to automate co-evolution between production and test code, identifying obsolete test cases to improve software quality and reduce maintenance overhead. TestART [44] enhances unit testing for LLMs by combining automated test generation with repair, using template-based fixes and coverage feedback to improve test accuracy.
@@ III-A Framework Overview
An overview of CoCoEvo is presented in Fig. 2. The evolutionary process begins with the random initialization of both the program population and the test case population, achieved through the random generating programs and test cases using LLMs. The co-evolution process follows a loop consisting of two alternating steps: program evolution and test case evolution, each of which includes a Generation-Evaluation-Selection sequence. During program evolution, the test case population is used to evaluate programs and calculate their fitness. Conversely, during test case evolution, the program population is utilized to evaluate test cases.
For programs, we introduce LLM-based crossover and LLM-based mutation operators for offspring generation. For test cases, we implement an LLM-based additional test case generation operator. Once the preset number of iterations is reached, the co-evolution loop terminates, and the program with the highest fitness score is returned as the final solution. Algorithm 1 provides the pseudo-code detailing the CoCoEvo process. Further implementation details for each algorithm step are elaborated in subsequent sections.
@@ III-B Program Evolution
This section provides a detailed description of strategies employed in program evolution.
Program fitness function. During evaluation, programs and test cases undergo cross-evaluation, as illustrated in Fig. 3. The outcomes are represented by a matrix :
We compute the confidence score of each program, adopting the approach in [12]. Programs that pass the same set of test cases are grouped into a set , while the corresponding test cases are grouped into a set . The confidence of program is then defined as
where represents the set containing program , and represents the associated test set. The confidence score evaluates the likelihood of a program being correct. The fitness of program is defined as
@@
Crossover Rate Scheduler To address the multimodal nature of the program search space and mitigate the risk of premature convergence, a dynamic scheduler is employed to regulate crossover and mutation rates. A higher mutation rate is applied in the early stages to promote exploration, while the mutation rate decreases, and the crossover rate increases in later stages to refine convergence.
Unlike traditional random mutation methods, the LLM-based program mutation operator leverages LLMs for code generation, which could retain key features of high-quality programs. This allows for a higher mutation rate without significant loss of desirable traits. We adopt the cosine annealing method [45] for scheduling crossover rate:
This formulation ensures that the crossover rate increases smoothly from to . The number of crossover operations and the number of mutation operations are defined as follows:
where is the program population size, and represent the number of program crossover and mutation operations in iteration round , respectively. When the initialization is completed, the iteration round is set to . After initialization, the iteration round starts at and increments by after a co-evolution loop.
Fig. 4 illustrates the variation in crossover rates over evolutionary rounds. This approach ensures a smooth transition between exploration (early stages) and convergence (later stages), enabling effective optimization of the solution space.

### v2

@@ Abstract
Large Language Models (LLMs) have shown remarkable performance in automated code generation. However, existing approaches often rely heavily on pre-defined test cases, which become impractical in scenarios where such cases are unavailable. While prior works explore filtering techniques between programs and test cases, they overlook the refinement of test cases. To address this limitation, we introduce CoCoEvo, a novel LLM-based co-evolution framework that simultaneously evolves programs and test cases. CoCoEvo eliminates the dependency on pre-defined test cases by generating both programs and test cases directly from natural language problem descriptions and function headers. The framework employs specialized evolutionary operators, including LLM-based crossover and mutation operators for program evolution, along with an additional test case generation operator for test case evolution. Additionally, we propose optimization strategies such as a crossover rate scheduler to balance exploration and convergence, and a multi-objective optimization method for test case selection. Experimental results on multiple state-of-the-art LLMs demonstrate that CoCoEvo surpasses existing methods, achieving state-of-the-art performance in automated code generation and testing. These results underscore the potential of co-evolutionary techniques in advancing the field of automated programming.
@@
Despite their impressive performance, LLMs do not always generate accurate or correct programs. To enhance the reliability of program generation, test cases are often utilized to evaluate code correctness and provide feedback. Existing approaches frequently rely on pre-defined test cases for evaluation. For instance, Sampling+Filtering [5] first generates a batch of candidate codes, evaluates them using pre-defined test cases, and selects the code with the highest score as the final answer. Other methods, such as Reflexion [6], Self-Repair [7], INTERVENOR [8], LDB [9], and PairCoder [10], utilize execution feedback on pre-defined test cases to refine or repair generated code. These methods generally assume that pre-defined test cases are accurate and can be fully trusted.
However, pre-defined test cases are not always available. In test-driven development (TDD) [11], test cases need to be defined before program implementation. In real-world software development, teams often rely solely on natural language requirements, as there are no pre-existing test cases or example programs. Furthermore, creating pre-defined test cases can be labor-intensive, requiring significant manual effort. In situations where LLMs must generate test cases independently, the effectiveness of existing methods remains uncertain.
Some approaches do not rely on pre-defined test cases. For instance, methods like CodeT [12] and MBR-Exec [13] generate both code and test cases using LLMs and filter the code based on the execution results of these test cases. Similarly, CodeCoT [14], which utilizes chain-of-thought prompting [15], enables LLMs to generate code and test cases simultaneously and evaluate and refine the code based on these cases. Additionally, AgentCoder [16] introduces a multi-agent system with a dedicated Tester agent to generate test cases for evaluating and repairing code. However, studies such as [12, 17] highlight that LLM-generated test cases can contain significant errors, potentially leading to unreliable evaluations and incorrect feedback. Moreover, many of these methods fail to verify the quality of the generated test cases. While approaches implement basic filtering mechanisms for test cases, other methods often use generated test cases directly for code evaluation without ensuring their correctness.
@@
To address these limitations, we propose CoCoEvo, an LLM-based co-evolution framework that enables the simultaneous evolution of programs and test cases to produce more accurate outcomes (illustrated in Fig. 1). In CoCoEvo, both programs and test cases are generated by LLMs based on natural language problem descriptions. No pre-defined programs, test cases, or templates are needed. The CoCoEvo framework consists of two alternating steps: program evolution and test case evolution. In program evolution, we introduce LLM-based crossover and mutation operators to generate program offspring. The current test case population is used to calculate program fitness. In test case evolution, we design an additional test case generation operator to produce test case offspring. Conversely, the current program population is used to calculate test case fitness. For programs, we implement a fitness calculation method inspired by CodeT [12], which assesses agreement between programs and test cases. For test cases, we adopt a multi-objective optimization approach to select those with high accuracy and strong discriminatory power, using the Pareto method [22]. Additionally, we introduce a dynamic crossover rate scheduling mechanism, enabling broader exploration during early evolution stages and faster convergence later.
To evaluate the effectiveness of CoCoEvo, we collected a LeetCode-Contest dataset comprising relatively new programming problems that are unlikely to have been included in the training data of the evaluated LLMs. We conducted experiments on four state-of-the-art models: GPT-4o-mini [1], Qwen2.5-Coder-32B [3], Llama-3.1-70B [2], and DeepSeek-V3 [4]. We compared with both basic code generation methods, repair-based methods, and agreement-based methods. The results demonstrate that CoCoEvo outperforms all existing methods. Additionally, the experiments revealed that methods relying on pre-defined test cases for feedback experience significant performance degradation when adapted to self-generated test cases. This highlights the limitations of such methods in scenarios lacking pre-defined test cases. Furthermore, we conducted in-depth comparative experiments and discussions, which validated the effectiveness of the optimization strategies employed in CoCoEvo.
@@
Optimized Evolution Techniques. For program evolution, we design a dynamic crossover rate scheduler that enhances exploration early on and accelerates convergence later. For test cases, we develop a multi-objective optimization method for fitness calculation and new population selection.
@@
Experimental Validation. We introduce the LeetCode-Contest dataset and conduct experiments using four state-of-the-art LLMs: GPT-4o-mini, Qwen2.5-Coder-32b, Llama-3.1-70B, and DeepSeek-V3. The experimental results demonstrate that CoCoEvo achieves state-of-the-art performance, surpassing existing methods. Additionally, we conduct comparative experiments to evaluate the utility of each module within the framework, further highlighting the effectiveness of our approach.
The remainder of this article is organized as follows. Section II discusses the related work. In Section III, the design details and implementation methods of the CoCoEvo are described. Sections IV and V present the experimental settings and discussions, respectively. Finally, Section VI concludes this paper and discusses the limitations and future work.
@@ II-A Enhancing Code Generation Accuracy with Test Case
Integrating test cases is now crucial for improving LLM-generated code. Key methodologies use them to filter, evaluate, and refine programs, enhancing performance and reliability.
Li et al. [5]’s Sampling+Filtering approach samples numerous candidate programs and scores them against a given test set. Chen et al. [12] use LLMs to co-generate code and tests, selecting superior programs via a dual agreement metric. Shi et al. [13] use test execution results to assess program similarity, guiding program selection.
Beyond evaluation, test-case results drive iterative repair and refinement of generated code. Reflexion [6] converts test failures into verbal feedback for program repair. Similarly, Olausson et al. [7] evaluate LLMs’ capabilities to repair programs using failed test feedback. Zhang et al. [23] allow LLMs to analyze test execution results to revise erroneous code. Zhong et al. [9] extend this idea by enabling step-by-step debugging. CodeCoT [14] uses chain-of-thought prompting [15] to sequentially generate code and tests, iteratively refining the code with test feedback. Zhang et al. [10] introduce a pair programming framework where test feedback informs repair strategies and planning.
@@ II-B Automated Test Case Generation
Early works, such as [24], leveraged search-based methods to generate test cases. More recently, research has shifted toward using LLMs for this task. Studies such as [25], [26], and [27] have explored the effectiveness of LLMs in automating program testing. Similarly, [28] investigates methods for enhancing existing test cases using LLMs. Others, like TOGA [29], TOGLL [30], and [31], generate test oracles from predefined templates. Meanwhile, [32] trained a deep reinforcement learning model for text-to-test case generation. The application of LLMs in software development extends to generating test cases for various scenarios. For instance, [33] employs GPT-4 to generate security tests for identifying vulnerabilities, while [27] uses LLMs to assist in generating unit tests for automatic software testing. Similarly, [34] demonstrates the use of LLMs to create test cases aimed at reproducing general software bugs. LLMs are also employed to improve the code accuracy through test case generation. CodeT [12], for example, utilizes LLMs with zero-shot prompts to directly generate test cases. Reflexion [6] introduces a Test Agent powered by LLMs for test case generation.
@@ II-C Evolution with LLMs for Code Generation
The integration of LLMs with evolutionary algorithms has been extensively explored to enhance code generation. Bradley et al. [35] introduced OpenELM, an open-source Python library that designs specialized evolutionary operators for code generation. Chen et al. [36] employed LLMs as adaptive operators for evolutionary neural architecture search to optimize network designs. Liu et al. [37] proposed using LLMs to evolve optimization algorithms, automating the processes of initialization, selection, and mutation. Liu et al. [38] introduced a framework integrating LLMs with evolutionary computation to co-evolve heuristic descriptions and executable code. Ma et al. [39] combined LLMs with evolutionary search to generate and refine RL reward functions via a reward-reflection mechanism. Meyerson et al. [40] used few-shot prompting to create a domain-agnostic crossover operator by concatenating parent solutions into one prompt. Similarly, Romera et al. [41] integrated LLMs with evolutionary algorithms to create context-aware program mutations. Ye et al. [42]’s ReEvo combines evolutionary search and LLM-driven reflections to generate heuristics via verbal optimization gradients. AutoTest [43] embeds LLM co-generation of code and test cases in an evolutionary loop for improved optimization.
Additionally, various LLM-based co-evolution methods have been proposed. For instance, Chi et al. [19] identifies outdated test cases by co-evolving production and test code. Ruan et al. [20] constructs test oracles from bug reports using predefined templates. Chi et al. [19] used LLMs with dynamic validation to automate production and test code co-evolution, identifying obsolete tests to enhance software quality and lower maintenance overhead. TestART [44] enhances LLM unit testing by automatically generating and repairing tests, using template-based fixes and coverage feedback to improve accuracy.
@@ III-A Framework Overview
An overview of CoCoEvo is presented in Fig. 2. The process begins by randomly initializing both the program and test-case populations, using LLMs to generate the initial programs and test cases. The co-evolution process follows a loop consisting of two alternating steps: program evolution and test case evolution, each of which includes a Generation-Evaluation-Selection sequence. During program evolution, the test case population is used to evaluate programs and calculate their fitness. Conversely, during test case evolution, the program population is utilized to evaluate test cases.
For programs, we introduce LLM-based crossover and mutation operators for offspring generation. For test cases, we implement an LLM-based additional test case generation operator. Once the preset number of iterations is reached, the co-evolution loop terminates, and the program with the highest fitness score is returned as the final answer. Algorithm 1 provides the pseudo-code detailing the CoCoEvo process. Further implementation details for each algorithm step are elaborated in subsequent sections.
@@ III-B Program Evolution
Program evaluation. During evaluation, programs and test cases are cross-evaluated, as illustrated in Fig. 3. The outcomes are represented by a matrix :
where and represent the ‑th program and the ‑th test case, respectively.
We design a Confidence metric to evaluate program fitness, drawing inspiration from CodeT [12]. Programs that pass the same set of test cases are grouped into a set , while the corresponding test cases are grouped into a set . The Confidence of program is defined as
where the size of and are denoted by and , respectively. This Confidence score evaluates the likelihood of a program being correct. The fitness of program is defined as
@@
Crossover Rate Scheduler. When leveraging LLMs to perform program crossover and mutation, the crossover operation typically synthesizes the strengths of disparate programs but may drive the population toward premature convergence, whereas the mutation operation often introduces novel features yet cannot guarantee evolution in a more desirable direction. Therefore, we propose a dynamic crossover rate scheduler that adaptively adjusts the number of crossover and mutation operations. We define the crossover rate as the proportion of offspring programs generated through the crossover operator. In the early stages, a lower crossover rate is applied to promote exploration, and the crossover rate increases in later stages to refine convergence. This enables extensive exploration of the search space during the early stages of evolution, while facilitating rapid convergence toward superior solutions in later generations.
Unlike traditional random mutation methods, the LLM-based program mutation operator leverages LLMs for code generation, which could retain key features of high-quality programs. This allows for a higher mutation rate without significant loss of desirable traits. Thus, we adopt the cosine annealing method [45] for scheduling crossover rate:
where denotes the crossover rate at the -th iteration, and represent the initial crossover rate and the final crossover rate, which are set to and , respectively. During the iteration process, starts at and ends at . In the formula, both and are decremented by , since in the first iteration () the program population is doing initialization and no crossover or mutation operations are performed.
This formulation ensures that the crossover rate increases smoothly from to . In each generation, offspring programs are produced. The number of crossover operations is determined by multiplying the crossover rate by and taking the floor of the result. The number of mutation operations is equal to minus the number of crossover operations, as calculated by the following equations:
where is the program population size, and represent the number of program crossover and mutation operations in the -th iteration, respectively.
Fig. 4 presents an example of the variation in crossover rate over iterations. Since when , the population is randomly initialized, the figure illustrates changes in the crossover rate from to , with the parameter set to . This approach ensures a smooth transition between exploration (early stages) and convergence (later stages), enabling effective optimization of the solution space.